OpenStudy (anonymous):

use the geometric mean to find the 7th term in a geomtric sequence if the 6th term is 75 and the 8th term is 48.

OpenStudy (cwrw238):

r^2 = 48/75 = 16/25 r = 4/5 7th term = 75 * 4/5 = 60

OpenStudy (anonymous):

@cwrw238 did exactly what almost everyone (including me) would do, but the question said "use geometric mean" so you can also get the answer of 60 by computing \[\sqrt{75\times 48}=60\]

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

- i must admit i forgot what geometric mean was..